Suppose children in two schools (A and B) received chicken pox vaccines at the beginning of the school term. An outbreak of chicken pox occurred in the area and both schools were monitored for chicken pox cases. Children at School A (n = 250) were followed for 12 months and those at School B (n=300) were followed for 9 months. Using person-time to compare the two populations would yield a. Incidence rates b. Cumulative incidence c. Incidence density d. Prevalence rates 7. Calculate the total number of person years in the above study
Answer #1

Answer:

In school 'A', no. of children is 250 and the time of observation is 12 months or 1 year

so the no. of person year for school 'A' = 250 x 12 = 3000 person months or 250 person years

In school 'B', no. of children is 300 and the time of observation is 9 months or 0.75 years

so the no. of person year for school 'B' = 300 x 09 = 2700 person months or 225 person years

Prevalence rates for for both group is 7.

Prevalence = (No. of Cases having the disease / Population size) x 1000

No. of cases having the disease for School 'A' = (7/250) x 1000 = 28

No. of cases having the disease for School 'B' = (7/300) x 1000 = 23.34

a. Incidence rate = (No. of new cases / Total no. of people or population at risk) x 1000

Incidence rate for School 'A' = (28 / 250) x 1000 = 112 per thousand

Incidence rate for School 'B' = (23.34 / 300) x 1000 = 77.8 per thousand

b. Cumulative Incidence = (No. of disease onsets / No. initially at risk) x 1000

For School 'A' = (28 / 250) x 1000 = 112 per thousand

For School 'B' = (23.34 / 300) x 1000 = 77.8 per thousand

112 + 77.8 = 189.8 per thousand

c. Incidence density = (No. of disease onsets / Sum of person - time at risk) x 1000

No. of person year for school 'A' = 250 person - years

No. of person year for school 'B' = 225 person - years

Incidence density = (28+23.34) / (250+225) x 1000

108.08 per thousand

d. Total no. of person years in study

250 person years + 225 person years = 475 person year